OpenStudy (anonymous):

A) Find the binomial probability P(x = 5), where n = 12 and p = 0.70. (B) Set up, without solving, the binomial probability P(x is at most 5) using probability notation. (C) How would you find the normal approximation to the binomial probability P(x = 5) in part A? Please show how you would calculate µ and σ in the formula for the normal approximation to the binomial, and show the final formula you would use without going through all the calculations.

OpenStudy (anonymous):

A.) P(x = 5) = C(12,5)(.3^5)(.7^7) = .158 b.) P(x<=5) = P(x=0) + P(x=1)+...+P(x=5). So, P(x=i) = C(12,i)(.3^i)(.7^(12-i)) C.) Use µ = .3(12)= 3.6, σ = sqrt(12(.3)(.7)) = 1.5874 If N is distributed with µ = 3.6, σ = 1.5874 Then, P(x=5) ~ P(N<5.5) - P(N <4.5).

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yw!